OpenStudy (anonymous):

Use the method of Lagrange multipliers to fi nd the maximum and minimum values of the function f(x;y) = xyz subject to the constraint x^2 + 2y^2 + 3z^2 = 6.

OpenStudy (anonymous):

yz=2xλ xz=4yλ xy=6zλ x^2 + 2y^2 + 3z^2 = 6. yz/2x=λ xz/4y=λ xy/6z=λ yz/2x=xz/4y 2y^2=x^2 xz/4y=xy/6z 3z^2=2y^2 2y^2+2y^2+2y^2=6 6y^2=6 y=+/-1 y=1: z=+/-sqrt(2/3); x=+/-sqrt(2) y=-1:z=+/-sqrt(2/3); x=+/-sqrt(2) critical points: (1,sqrt(2/3),sqrt(2))--->a (-1,-sqrt(2/3),-sqrt(2)--->b f(a)=sqrt(2/3)*sqrt(2)..............maximum f(b)=-sqrt(2/3)*sqrt(2)............minimum i apologize if i made any mistakes i haven't done these since last term